# Tailoring anisotropy in kirigami metamaterial skins with pop-up folding hinges

### v1

@@ 5.2 Finite element model
We model kirigami sheets in Abaqus with S4R5 shell elements; the connections between pop-ups and octagonal panels are modeled using connector elements. The dimensions are chosen to be the same as those of the physical samples. Panels are modeled with a linear elastic material model having Young’s modulus of 2.2 GPa and Poisson’s ratio of 0.38, taken from measurements mentioned in Ref. [55]. The joint parameters are chosen to allow rotations about in-plane axes while restricting rotations about the out-of-plane axis; unless otherwise specified. The torsional stiffness about the in-plane and out-of-plane axes are chosen to be $k_{tx}=k_{ty}=10^{-6}$ and $k_{tz}=10^{3}$, respectively. Preliminary simulations suggest that a minimum of three hinges are necessary to properly connect each pop-up panel to its corresponding octagonal panel; since the smallest fold length in our representative geometries is $0.2\,L_{x}$ with $L_{x}=4.5\,$cm, we place a hinge every 3 mm of fold line. Note that choosing the right number of connector elements per fold would have been much more important if we were to compare simulations and experiments in terms of force-deformation responses [56] — something that we do not do here.
We perform two sets of planar simulations, depending on their intended use. We first build a unit cell model with periodic boundary conditions to simulate the behavior and kinematics of infinite sheets. Details on the implementation of periodic boundary conditions are given in B. In these simulations, panels are not constrained to remain in-plane, as these constraints cause convergence issues; we will see later that the panels naturally remain almost completely planar. We use these periodic models to compare to the analytical and experimental results of a single cell, thus validating whether single cells are representative of periodic specimens.
We also build a single cell model with boundary conditions akin to those previously introduced in Fig. 3B, including out-of-plane displacement restraints for all panel nodes, and subjected to loading along the horizontal direction at the top-most node of the bottom-right folding hinge. We mainly use this model to compute stress-strain curves and to compare them to those of the periodic model to get an idea on the mechanics of these sheets. In both sets of simulations, since the pop-up of hinges is a buckling problem, we perform the analysis in two steps: first, we perform a linear buckling analysis and identify critical buckling loads and modes of interest, corresponding to the desired kinematics. These buckling modes are then used as imperfections in post-buckling investigations. Post buckling analyses are performed in force-control mode with a quasi-static dynamic implicit method, with a sufficiently large time period chosen to minimize the effects of inertial forces.
Fig. 9: Response of representative patterns. (A) Response of a kirigami with mechanism-like behavior and with $\bar{w}_{x}=\bar{w}_{y}=0$, ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$, $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$. (B) Response of a kirigami with mechanism-like behavior and with $\bar{w}_{x}=0.3$, $\bar{w}_{y}=0$, ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$, $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$. (C) Response of a kirigami with mechanism-like behavior and with $\bar{w}_{x}=\bar{w}_{y}=0$, ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$, $\lambda_{1}=\lambda_{3}=\lambda_{4}=0.39$, $\lambda_{2}=1.15$. (D) Response of a kirigami with mechanism-like behavior and with $\bar{w}_{x}=\bar{w}_{y}=0$, ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=0.2$, ¯$\bar{\ell}_{2}=\bar{\ell}_{4}=0.7$, $\lambda_{1}=\lambda_{3}=\lambda_{4}=0.45$, $\lambda_{2}=0.85$. (E) Response of a kirigami without mechanism and with $\bar{w}_{x}=\bar{w}_{y}=0$, ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.2$, ¯$\bar{\ell}_{2}=0.4$, $\lambda_{1}=\lambda_{3}=\lambda_{4}=\pi/6$, $\lambda_{2}=\pi/4$. Throughout the figure, ($\square$-1) subfigures show the geometry, ($\square$-2) show a numerically-simulated snapshot of the deformed structure, ($\square$-3) and ($\square$-4) show a comparison between analytical, numerical and experimental results for the strains and pop-up heights, respectively, as a function of the actuation angle $\theta$, while ($\square$-5) show the stress-strain numerical response. Experimental results represent the mean and standard deviation of three measurements.
5.3 Results
We begin the analysis of the patterns listed in Table 1 by studying the auxetic one with symmetric pop-ups shown in Fig. 9A-1 (akin to the one by Tang et al. [37]). Fig. 9A-2 illustrates a snapshot of the FE simulation of the same unit, which allows to see the mesh used for its analysis. Quantitative results are shown in Figs. 9A-3,A-4. From the plot of strains vs. $\theta$ in Fig. 9A-3, we can see that numerical simulations and analytical calculations overlap perfectly, and that the experimental data points follow a very similar up-down trend. The pop-up height plots in Fig. 9A-4 also show a remarkable consistency with experiments. While there are small discrepancies with the numerical simulations, these are due to the fact that, in these simulations with periodic boundary conditions, we do not constrain the out-of-plane displacement of the octagonal panels; thus, small out-of-plane deflections that can alter the pop-up heights are expected. We noticed that these discrepancies get smaller with stiffer sheets, that are less prone to panel bending. We now turn to the non-auxetic pattern with symmetric pop-ups of Figs. 9B-1,B-2. The quantitative results in Figs. 9B-3,B-4 show again good consistency between analytics and experiments, with the numerics following similar but not identical trends. The fact that pop-ups in this pattern behave in the same way as those of the auxetic pattern in Fig. 9A-1, despite significantly different global morphing behaviors, highlights that pop-ups are relatively independent of planar kinematics, as we claimed multiple times in previous sections.
We now move to the analysis of patterns with asymmetric pop-ups, starting from the one in Fig. 9C-1, whose FE model at some level of deformation is shown in Fig. 9C-2. The strain evolution of Fig. 9C-3 shows that, in this case, numerics and analytics underpredict the extension achieved by the experimental sample, especially for the highest-strain data point; this is attributed to the fact that this data point represents a configuration where some pop-ups are nearly fully folded, and that full folding is not easy to achieve in our experimental samples due to an inevitably non-zero folding radius. Yet, the pop-up height plots in Fig. 9C-4 show again remarkable consistency for all pop-ups. Similar considerations can be made for the non-auxetic pattern with asymmetric pop-ups shown and described in Figs. 9D-1–D-4; in this case, however, analytical, numerical, and experimental results are more consistent than in the auxetic, asymmetric case. Finally, we resort to numerical simulations alone to study the pattern in Figs. 9E-1,E-2; owing to the fact that ${\ell}_{2}$ is different from all other hinge lengths, this pattern does not feature a mechanism and therefore cannot be studied with our analytical kinematics tools, and its morphing attributes cannot be easily quantified in experiments. The strain plot in Fig. 9E-3 shows that this pattern has mild auxeticity, but less than the similar pattern with mechanism of Fig. 9C-1. This is due to the fact that, as shown in Fig. 9E-2, the panels undergo significant bending. The absence of mechanisms also causes the pop-ups to be all different from each other, as shown in Fig. 9E-4. Interestingly, one of the pop-ups has negative height, owing to the fact that it borders a panel that bends significantly.
Some key aspects of the mechanics of these kirigami patterns, studied via numerical simulations alone, are summarized in Fig. 10.
Fig. 10: Preliminary information on the mechanics of planar sheets. All graphs are stress-strain curves. (A) Mechanism loss due to increasing torsional stiffness about in-plane axes in the unit of Fig. 9A-1. The torsional stiffness about the out-of-plane $z$ axis is set to $10^{3}$. Snapshots of the deformation of that same pattern, with (B) $k_{tx}=k_{ty}=10^{-6}$ or $10^{-4}$, and (C) $k_{tx}=k_{ty}=1$. The mesh here is identical to that of Fig. 9A-2. Comparison between stress-strain curves for all patterns described in Fig. 9, with (D) periodic boundary conditions, or (E) referring to a single isolated cell.
First, we analyze the effects of the in-plane torsional stiffness $k_{tx}$ (equal to $k_{ty}$) on the response of the pattern of Fig. 9A-1, with periodic boundary conditions. The stress-strain curves for three different stiffness values are shown in Fig. 10A, which includes a full-scale figure (right) and a zoom-in of it (left). Stress here is intended as the total applied force on the unit cell, normalized by the unit’s cross sectional area ($2L_{x}t$). We can see that choosing a small $k_{tx}$ ($10^{-6}$ or $10^{-4}$) yields curves where the force rapidly increases and then plateaus as in Euler buckling problems, with the plateau here corresponding to the mechanism mode of deformation. Note that the curve for $10^{-6}$ curls back towards zero strain as boundary conditions in periodic patterns (essentially constraining regions along the boundary to behave the same way, without imposing displacements or forces on them) allow them to do so. While we can see that the energy needed to actuate these patterns varies with the torsional stiffness, the mechanism parts of the curve are very similar, and the kinematics identical. For this reason, we choose $k_{tx}=k_{ty}=10^{-6}$ throughout this paper, as this value guarantees that the mechanism is present. Choosing larger values, e.g., $k_{tx}=k_{ty}=1$, yield instead a much stiffer response where the curve does not feature a plateau, as shown in the right panel of Fig. 10A. The deformed unit cells in Fig. 10B,C, corresponding to low and high $k_{tx}$, respectively, show indeed that the mechanism is lost when $k_{tx}=k_{ty}=1$ as highlighted by the overall bending of the unit cell and by the absence of pop-ups. As a side note, this is the type of mechanics response we would expect for specimens like the ones we fabricated here, owing to the high torsional stiffness of their hinges.
We conclude this analysis by comparing stress-strain curves for all patterns in Table 1. Fig. 10D shows such comparison for simulations with periodic boundary conditions. First, we observe that, as expected, the pattern without a mechanism (Fig. 9E-1) is much stiffer than the others. We can also see that out-of-plane buckling of patterns in Figs. 9A-1,B-1,C-1 involve similar forces, while it takes higher forces to pop-up the pattern in Fig. 9D-1. The mechanism regions have different slopes, but it must be kept in mind that we are only plotting strains up until $\varepsilon_{x}\approx 0.07$ and that the pattern in Fig. 9D-1 can reach much higher strains than that; thus, its mechanism portion of the deformation curve is bound to increase at larger strains. For completeness, we also report the stress-strain curves obtained from simulations on isolated unit cells. In this case, the different patterns display an even larger variability in terms of response. Further interpretations on these results could be given following energetic arguments similar to those in Ref. [37] and based on torsional springs, lengths of folding lines and angles of rotations between pop-ups and panels; we leave these types of mechanistic considerations for future work, where the considerations could be experientally validated.
There are a few important take-home messages from the analysis of these planar patterns:
  • 1.
    the consistency between analytics, numerics and experiments indicates the reliability of our modeling techniques; in particular, the analytical formulas derived here allow for a very quick and comprehensive analysis of the kinematics of these patterns, while the numerical simulations allow to only characterize the unit cell’s behavior for increasing $\varepsilon_{x}$, and at a much higher computational cost;
  • 2.
    yet, numerical simulations and the mechanics considerations they enable are still very important when designing such systems, and indicated by the fact that the force needed to actuate these patterns also varies significantly as a function of the design parameters;
  • 3.
    changing a few design parameters allows to completely alter the behavior of kirigami with folding hinges, highlighting the tailorability of these designs.
Fig. 11: Numerical simulations of tubular kirigami with asymmetric pop-ups. (A-1) Tubular arrangement of the unit cell of Fig. 9C-1. Deformed unit cells at axial strain values: (A-2) $\varepsilon_{x}=0.0055$, (A-3) $\varepsilon_{x}=0.0218$, (A-4) $\varepsilon_{x}=0.0572$. (B-1) Tubular arrangement of the unit cell of Fig. 9D-1. Deformed unit cells at axial strain values: (B-2) $\varepsilon_{x}=0.0743$, (A-3) $\varepsilon_{x}=0.144$, (A-4) $\varepsilon_{x}=0.167$.
6 From sheets to tubular kirigami shells
We conclude this work by briefly illustrating how the proposed geometries behave when arranged into tubular shapes. We do so using numerical simulations alone. The key assumption here is that the analyzed architected shell is stress free in its undeformed configuration; this would not be the case if this shell were to be fabricated by wrapping an initially-flat sheet around itself to create a tube. The configurations we study are those that present asymmetric pop-ups: the auxetic one of Fig. 9C-1, and the non-auxetic one of Fig. 9D-1. Our tubes feature 32 unit cells in the circumferential direction, as shown by the tubular slices in Figs. 11A-1 and B-1 for the auxetic and non-auxetic patterns, respectively. As done in other works [40], we model our tubular kirigami using single unit cells and periodic boundary conditions. Details on the implementation of periodic boundary conditions for tubular patterns are also given in B. The solution procedures and elements used for these simulations are identical to those for the planar simulations, except that we use displacement-controlled instead of force-controlled loading.
The deformation of the auxetic tube for different values of axial elongation are shown in Figs. 11A-2–A-4, where the colormap indicates radial displacement. As expected for an auxetic geometry, we can see that the tube increases its radius in response to an axial extension. We can also see that pop-ups raise asymmetrically as in the planar analog. If this tube were to be wrapped around an actuator to create a crawling or burrowing robot, it is expected that this robot would move more easily along the $+x$ direction than along the $-x$ direction. Thus, this pattern is a great candidate to create a robotic skin that is auxetic and has anisotropic friction. Similarly, in Figs. 11B-2–B-4, we show the deformation of the non-auxetic pattern which, as expected, produces a tube that shrinks in response to axial elongation. We can see that, in this case, the pop-ups with backbone aligned along the axial direction are very pronounced. We can also see that one of those pop-ups is more pronounced than the other, thus making this skin a good candidate to display non-auxeticity and anisotropic friction.
7 Conclusions
In this work, we performed a comprehensive study on the morphing attributes of kirigami skins with folding hinges by means of analytical kinematics studies, numerical simulations and experiments, and showed how to properly parameterize these patterns to obtain sheets with vastly-different behaviors. One significant finding is that such kirigami display global mechanism modes of deformation characterized by counter-rotating panels and hinge pop-ups as long as hinge lengths are all identical or symmetric about the horizontal and vertical axes of the unit cell; in other words, the quadrilaterals joining the rotation points of four adjacent panels need to be parallelograms. This finding rests on our conjecture on the hierarchy of mechanisms, i.e., that the pop-up hinges do not affect the global mechanism — which we do not prove formally but instead substantiate via extensive numerical and experimental verifications. The other significant finding is that the folding hinge angles can all be tuned independently, within limits that avoid panel overlaps. As a result of these findings, this kirigami platform allows to design patterns with vastly-different global morphing behavior, from auxetic to non-auxetic with various expansion capacity along different directions, and local morphing behavior, with symmetric or asymmetric pop-ups. As such, it has potential to be used as skins in the context of crawling and burrowing robots, or in other applications where asymmetric morphing or anisotropic friction are desirable.
A natural follow up study could involve creating a crawling or burrowing robot prototype, evaluating its anisotropic friction, and further refining the search of the design space based on an improved understanding of what matters the most to maximize this anisotropy. One challenge to face in application-oriented studies is related to the fabricability of these patterns, which feature large stress concentrations at the points of intersection of various panels. This could be alleviated by fabricating panels and folding hinges as separate entities, and connecting them via rivets or other joints. Another challenge to overcome is related to the overall “floppiness” of these patterns; in fact, unless properly constrained, it is not certain that these patterns would follow the desired mechanism mode of deformation with counter-rotating yet co-planar panels. This challenge could be overcome by constraining some points of the panels to follow a stretchable substrate. Furthermore, directionally biasing the folding hinges might be needed to make sure that they pop-up towards the desired side of the sheet.
More fundamental extensions of this work, following analog studies for other kirigami systems [37, 49, 40], could be related to: i) introducing asymmetric folding hinges in other planar kirigami patterns such as the kagome one; ii) studying the complete out-of-plane kinematics of these patterns, without constraining the large panels to remain in-plane; or iii) investigating how pop-ups propagate in these patterns and how to tune unit cell parameters across a tube to tailor the pop-up sequence.
Acknowledgments
P.C. and A.K. acknowledge the support of the National Science Foundation (CMMI-2228271 and CMMI-2228272). A.W. acknowledges the support of Stony Brook University through a New York Space Grant research award. We thank Ahmad Rafsanjani, now at the University of Southern Denmark, for sharing his Abaqus python codes, from which we learned how to apply periodic boundary conditions and simulate tubular patterns.
Author Contributions
Appendix A Pellegrino-Calladine analysis details
In this appendix, we detail the matrix analysis used in Section 3 to determine which sets of parameters yield mechanisms.
Once the truss analog to our unit created, we compile its equilibrium matrix $\mathbf{A}\in\mathbb{R}^{(nj)\times b}$, that relates bar tensions $\mathbf{t}\in\mathbb{R}^{b}$ to nodal forces $\mathbf{f}\in\mathbb{R}^{nj}$ as follows:
$\mathbf{A}\mathbf{t}=\mathbf{f}.$ (11)
This matrix can be computed by taking the equilibrium of forces about each joint of the structure and assembling all equations in matrix form. Similarly, we can write a statement of kinematics as the relationship between bar elongations $\mathbf{e}\in\mathbb{R}^{b}$ and nodal displacements $\mathbf{d}\in\mathbb{R}^{nj}$, via the kinematic matrix $\mathbf{B}\in\mathbb{R}^{b\times(nj)}$:
$\mathbf{B}\mathbf{d}=\mathbf{e}.$ (12)
Equilibrium implies that $\mathbf{B}=\mathbf{A}^{\text{T}}$. This statement can be justified by putting together Eqs.(11) and (12) with a generic material model $\mathbf{t}=\mathbf{C}\mathbf{e}$, where $\mathbf{C}\in\mathbb{R}^{b\times b}$ is the matrix of material constants; the result is that $(\mathbf{A}\mathbf{C}\mathbf{B})\mathbf{d}=\mathbf{f}$, where $\mathbf{K}=(\mathbf{A}\mathbf{C}\mathbf{B})$ is the stiffness matrix; energetic arguments require $\mathbf{K}$ to be symmetric, which in turn requires that $\mathbf{B}=\mathbf{A}^{\text{T}}$.
The mechanisms of the structure can be found by computing the left null space of matrix $\mathbf{A}$ (i.e., the null space of $\mathbf{B}$); each column of this subspace is a set of displacements corresponding to a mechanism, while its number of linearly independent columns is the number of mechanisms $m$. Computing the mechanisms alone is not enough if we need information on their nature. In fact, in truss structures, we can have redundancies called states of self stress. Such states of self stress can stiffen mechanisms, making them infinitesimal in nature and not capable of expressing over finite deformations. To compute the states of self stress of a truss, we look at the null space of $\mathbf{A}$. Each row of it represents a set of bar tensions in equilibrium without external forces — a state of self stress; the number of linearly independent rows in this space is the number of states of self stress $s$. To verify whether states of self stress stiffen mechanisms, we compute the product forces needed to displace the structure like a mechanism with bar tensions coming from a state of self stress. These product forces are then used to augment the column space of $\mathbf{A}$, which represents the forces that the system can carry. We then study the rank of this augmented column space: if full, the state of self stress stiffens the considered mechanism; if not full, the mechanism is not affected by the state of self stress. This procedure is repeated for all combinations of mechanisms and states of self stress. If a mechanism is unaffected by all states of self stress, it is finite in nature.
Appendix B Periodic boundary conditions
Here, we provide further details on the implementation of periodic boundary conditions (BCs) within our Abaqus models for both planar and tubular kirigami.
B.1 Periodic BCs for planar kirigami
@@
The main challenge in implementing periodic boundary conditions for planar sheets is to correctly select boundary conditions to ensure stability and the application of the proper constraints on the unit cell. To do so, we basically set displacements of opposite fold lines to be equal and opposite (Fig. 2C is useful to understand which fold lines need to be related to which in a prototypical pattern); this is done through two dummy nodes, i.e., reference points not connected to the cell but whose displacement is featured in the equations coupling the displacements of related fold lines. The first dummy node imposes essential boundary conditions and deformations from fold lines on the top edge to those on the bottom edge, while the second dummy node incorporates both essential and natural boundary conditions, aligning the right edge of the sheet with the left edge. Loading is then applied along the horizontal direction at the dummy node.
B.2 Periodic BCs for tubular kirigami
@@
Periodic boundary conditions for tubular patterns require a slightly different implementation with respect to those for planar kirigami. In particular, we here pay special attention to imposing rotational constraints (for rotations along the axis of the tube) instead of displacement constraints between the pop-ups at the top and bottom of the cell, in order to maintain the overall symmetry and structural integrity of the system. In addition to the applied periodic boundary conditions, further essential boundary conditions are utilized to help the cylinders keep their axes constant and their cross section circular during deformation. To this aim, rotation of the octagonal panels in the axial direction of the unit cells are set to be zero.

### v2

@@ E.1 General model information
We model kirigami sheets in Abaqus with S4R5 shell elements; the connections between pop-ups and octagonal panels are modeled using connector elements [78]. The dimensions are chosen to be the same as those of the physical samples, with $L_{x}=L_{y}=2.5$ cm. Panels are modeled with a linear elastic material model having Young’s modulus of 2.2 GPa and Poisson’s ratio of 0.38, taken from measurements mentioned in Ref. [77]. The joint parameters are chosen to allow rotations about in-plane axes while restricting rotations about the out-of-plane axis; unless otherwise specified. The torsional stiffness about the in-plane and out-of-plane axes are chosen to be $k_{tx}=k_{ty}=10^{-6}$ and $k_{tz}=10^{3}$, respectively. Preliminary simulations suggest that a minimum of three hinges are necessary to properly connect each pop-up panel to its corresponding octagonal panel; thus, we place three connector elements along each fold line. As explained in the main text, this simplification is bound to create discrepancies between models and experiments.
E.2 Unit cell models
@@
We first build unit cell models with periodic boundary conditions to simulate the behavior and kinematics of infinite sheets. The main challenge in implementing periodic boundary conditions for planar sheets is to correctly select boundary conditions to ensure stability and the application of the proper constraints on the unit cell. To do so, we basically set displacements of opposite fold lines to be equal and opposite (Fig. 2(c) is useful to understand which fold lines need to be related to which in a prototypical pattern); this is done through two dummy nodes, i.e., reference points not connected to the cell but whose displacement is featured in the equations coupling the displacements of related fold lines. The first dummy node imposes essential boundary conditions and deformations from fold lines on the top edge to those on the bottom edge, while the second dummy node incorporates both essential and natural boundary conditions, aligning the right edge of the sheet with the left edge. Loading is then applied along the horizontal direction at the dummy node.
In these unit cell simulations, panels are not constrained to remain in-plane, as these constraints cause convergence issues; it is observed that the panels can naturally remain almost completely planar. We use these periodic models to compare to the analytical and experimental results of a single cell, thus validating whether single cells are representative of periodic specimens. Since the pop-up of hinges is a buckling problem, we perform the analysis in two steps: first, we perform a linear buckling analysis and identify critical buckling loads and modes of interest, corresponding to the desired kinematics. These buckling modes are then used as imperfections in post-buckling investigations. Post-buckling analyses are performed in force-control mode with a quasi-static dynamic implicit method, with a sufficiently large time period chosen to minimize the effects of inertial forces.
E.3 Tubular models
We also built models for tubes with octagonal cross sections and stiffened panels. Exploiting the structural and loading symmetry observed in experiments, we only modeled a quarter of the geometry, with periodic boundary conditions enforcing the symmetry. We used a cylindrical coordinate system to define these conditions, which, similar to previous models, involved displacement constraints only. We then applied boundary conditions akin to those in the tensile test, reflecting the rpesence of end caps. Consequently, we used a single dummy node to define the displacement constraints along the symmetry lines, and we applied the tensile load directly to the model’s nodes.
@@ E.3 Tubular models
Implementing boundary conditions for cylindrical and octagonal tubular patterns necessitates a distinct approach compared to planar kirigami. Notably, rotational constraints about the tube’s axis are prioritized over displacement constraints between the top and bottom pop-ups of the unit cell. This ensures the preservation of overall symmetry and structural integrity. To further maintain a constant tube axis and tubular cross-section during deformation, we apply essential boundary conditions, specifically enforcing zero axial rotation of the octagonal panels within the unit cells.
E.4 On the effects of the hinge stiffness
Fig. 12: Effects of hinge stiffness on the mechanics of single units. All graphs are stress-strain curves. (a) Mechanism loss due to increasing torsional stiffness about in-plane axes in the SN unit. The torsional stiffness about the out-of-plane $z$ axis is set to $10^{3}$. Snapshots of the deformation of that same pattern, with (b) $k_{tx}=k_{ty}=10^{-6}$ or $10^{-4}$, and (c) $k_{tx}=k_{ty}=1$.
In our work, as previously mentioned, we choose specific values for the torsional hinge stiffness. Our choice is explained in the following. To understand the effects of the in-plane torsional stiffness $k_{tx}$ (equal to $k_{ty}$) on the response of the pattern of Fig. 5(a-1), with periodic boundary conditions. The stress-strain curves for three different stiffness values are shown in Fig. 12(a), which includes a full-scale figure (right) and a zoom-in of it (left). Stress here is intended as the total applied force on the unit cell, normalized by the unit’s cross sectional area ($2L_{x}t$). We can see that choosing a small $k_{tx}$ ($10^{-6}$ or $10^{-4}$) yields curves where the force rapidly increases and then plateaus as in Euler buckling problems, with the plateau here corresponding to the mechanism mode of deformation. Note that the curve for $10^{-6}$ curls back towards zero strain as boundary conditions in periodic patterns (essentially constraining regions along the boundary to behave the same way, without imposing displacements or forces on them) allow them to do so. While we can see that the energy needed to actuate these patterns varies with the torsional stiffness, the mechanism parts of the curve are very similar in slope, and the kinematics identical. For this reason, we choose $k_{tx}=k_{ty}=10^{-6}$ throughout this paper, as this value guarantees that the mechanism is present. Choosing larger values, e.g., $k_{tx}=k_{ty}=1$, yield instead a much stiffer response where the curve does not feature a plateau, as shown in the right panel of Fig. 12(a). The deformed unit cells in Fig. 12(b) and (c), corresponding to low and high $k_{tx}$, respectively, show indeed that the mechanism is lost when $k_{tx}=k_{ty}=1$ as highlighted by the overall bending of the unit cell and by the absence of pop-ups.